OpenStudy (trwatkins1):

What is the value of –1.9 – q when q = 6.7?

OpenStudy (trwatkins1):

@YoungStudier

OpenStudy (youngstudier):

What do you think?

OpenStudy (trwatkins1):

4.3 (estimate)

OpenStudy (youngstudier):

Urm, you're half way right. XD

OpenStudy (trwatkins1):

yea thats why i put estimate

OpenStudy (youngstudier):

No as in multiply that by 2 and add a negative sign, half way right. XD

OpenStudy (trwatkins1):

4.-6

OpenStudy (trwatkins1):

im su confused

OpenStudy (youngstudier):

-1.9 - 6.7 = -1.9 + (-6.7) = -8.6

OpenStudy (trwatkins1):

oh

OpenStudy (youngstudier):

:D

OpenStudy (trwatkins1):

XD i thought the .3 was wrong

OpenStudy (youngstudier):

Urm, nah

OpenStudy (youngstudier):

:D

OpenStudy (youngstudier):

LOL, I'm 3 medals short of being able to get human calculator